Hello
350 is what percent of 1250?
Thank you

Found 2 solutions by solver91311, jim_thompson5910:
Answer by solver91311(24713)



Divide 350 by 1250. Multiply the answer by 100.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by jim_thompson5910(35256)

Let x = unknown percent

(Part)/(Whole) = (Percent)/100

(350)/(1250) = x/100

0.28 = x/100

100*0.28 = x

28 = x

x = 28

Answer: 350 is 28% of 1250